Please help me solve this.
2x+y=-10
3x+2y=-14

Answer by stanbon(75887)

Please help me solve this.
2x+y=-10
3x+2y=-14
---------------------
Let's say you choose to eliminate the y-terms.
--------
Multiply the 1st equation by 2:
4x+2y = -20
3x+2y = -14
-------------------
Subtract and solve for "x":
x = -6
---
Solve for "y":
2x+y = -10
2*-6 + y = -10
-12 + y = -10
y = 2
--------------------
Cheers,
Stan H.